### 29 Feb 2024 v1

# The 6th Affective Behavior Analysis in-the-wild (ABAW) Competition

Abstract
This paper describes the 6th Affective Behavior Analysis in-the-wild (ABAW) Competition, which is part of the respective Workshop held in conjunction with IEEE CVPR 2024. The 6th ABAW Competition addresses contemporary challenges in understanding human emotions and behaviors, crucial for the development of human-centered technologies. In more detail, the Competition focuses on affect related benchmarking tasks and comprises of five sub-challenges: i) Valence-Arousal Estimation (the target is to estimate two continuous affect dimensions, valence and arousal), ii) Expression Recognition (the target is to recognise between the mutually exclusive classes of the 7 basic expressions and ’other’), iii) Action Unit Detection (the target is to detect 12 action units), iv) Compound Expression Recognition (the target is to recognise between the 7 mutually exclusive compound expression classes), and v) Emotional Mimicry Intensity Estimation (the target is to estimate six continuous emotion dimensions). In the paper, we present these Challenges, describe their respective datasets and challenge protocols (we outline the evaluation metrics) and present the baseline systems as well as their obtained performance. More information for the Competition can be found in: https://affective-behavior-analysis-in-the-wild.github.io/6th.
1 Introduction
The 6th Affective Behavior Analysis in-the-wild (ABAW) Workshop and Competition continues its tradition of fostering interdisciplinary collaboration by bringing together experts from various fields including academia, industry, and government. This workshop, held in conjunction with IEEE Computer Vision and Pattern Recognition Conference (CVPR) 2024, aims to delve into the analysis of affective behavior in real-world settings, a critical aspect for the development of human-centered technologies such as HCI systems and intelligent digital assistants. By understanding human emotions and behaviors, machines can better engage with users irrespective of contextual factors like age, gender, or social background, thereby enhancing trust and interaction in real-life scenarios.
The ABAW Competition, an integral part of the workshop, is split into five Challenges.
The first Challenge is the Valence-Arousal (VA) Estimation one; the target of this Challenge is to estimate the two continuous affect dimensions of valence and arousal in each frame of the utilized Challenge corpora. Valence characterises an affective state on a continuous scale from positive to negative (in other words from -1 to 1). Arousal characterises an affective state on a continuous scale from active to passive (in other words from -1 to 1).
Only uni-task solutions will be accepted for this Challenge. Teams are allowed to use any -publicly or not- available pre-trained model (as long as it has not been pre-trained on the utilized in this Challenge corpora, Aff-Wild2). The pre-trained model can be pre-trained on any task (e.g., VA estimation, Expression Recognition, AU detection, Face Recognition). However when the teams are refining the model and developing the methodology they should not use any other annotations (expressions or AUs): the methodology should be purely uni-task, using only the VA annotations. This means that teams are allowed to use other databases’ VA annotations, or generated/synthetic data, or any affine transformations, or in general data augmentation techniques for increasing the size of the training dataset.
For this Challenge, an augmented version of the Aff-Wild2 [105, 44, 47, 37, 51, 40, 39, 54, 30, 31, 48, 59] is used. This corpora is audiovisual (A/V), in-the-wild and in total consists of 594 videos of around 3M frames of 584 subjects.
The second Challenge is the Expression (Expr) Recognition one; the target of this Challenge is to recognise between eight mutually exclusive classes in each frame of the utilized Challenge corpora; these classes are the 6 basic expressions (i.e., anger, disgust, fear, happiness, sadness and surprise), the neutral state and a category ’other’ that denotes affective states that are not included in the neutral state or in the 6 basic expressions.
Only uni-task solutions will be accepted for this Challenge. Teams are allowed to use any -publicly or not- available pre-trained model (as long as it has not been pre-trained on the utilized in this Challenge corpora, Aff-Wild2). The pre-trained model can be pre-trained on any task (e.g., VA estimation, Expression Recognition, AU detection, Face Recognition). However when the teams are refining the model and developing the methodology, they should not use any other annotations (VA or AUs): the methodology should be purely uni-task, using only the Expr annotations. This means that teams are allowed to use other databases’ Expr annotations, or generated/synthetic data (e.g. the data provided in the ECCV 2022 run of the ABAW Challenge [29]), or any affine transformations, or in general data augmentation techniques (e.g., [78, 60]) for increasing the size of the training dataset.
For this Challenge, the Aff-Wild2 is used. This corpora is audiovisual (A/V), in-the-wild and in total consists of 548 videos of around 2.7M frames.
The third Challenge is the Action Unit (AU) Detection one; the target of this Challenge is to detect which of the 12 Action Units are activated in each frame of the utilized Challenge corpora. Action Units refer to a set of facial muscle movements or configurations. The action units that have been selected for the purposes of this Challenge are the following: AU1, AU2, AU4, AU6, AU7, AU10, AU12, AU15, AU23, AU24, AU25 and AU26.
Only uni-task solutions will be accepted for this Challenge. Teams are allowed to use any -publicly or not- available pre-trained model (as long as it has not been pre-trained on the utilized in this Challenge corpora, Aff-Wild2). The pre-trained model can be pre-trained on any task (e.g., VA estimation, Expression Classification, AU detection, Face Recognition). However when the teams are refining the model and developing the methodology, they should not use any other annotations (VA or Expr): the methodology should be purely uni-task, using only the AU annotations. This means that teams are allowed to use other databases’ AU annotations, or generated/synthetic data, or any affine transformations, or in general data augmentation techniques (e.g., [78, 60]) for increasing the size of the training dataset.
For this Challenge, the Aff-Wild2 is used. This corpora is audiovisual (A/V), in-the-wild and in total consists of 542 videos of around 2.7M frames.
The fourth Challenge is the Compound Expression (CE) Recognition one; the target of this Challenge is to recognise between the 7 mutually exclusive classes in each frame of the utilized Challenge corpora. These classes are the following compound expressions: Fearfully Surprised, Happily Surprised, Sadly Surprised, Disgustedly Surprised, Angrily Surprised, Sadly Fearful and Sadly Angry.
Teams are allowed to use any -publicly or not- available pre-trained model and any -publicly or not- available database (that contains any annotations, e.g. VA, basic or compound expressions, AUs).
For this Challenge, a part of C-EXPR-DB [32] database will be used, which consists of 56 videos in total. C-EXPR-DB is an audiovisual (A/V) in-the-wild database and in total consists of 400 videos of around 200K frames.
The sixth ABAW Competition, which is part of the respective Workshop to be held in conjunction with the IEEE Computer Vision and Pattern Recognition Conference (CVPR) 2024 is a continuation of the successful series of ABAW Competitions held in conjunction with IEEE CVPR 2023, ECCV 2022, IEEE CVPR 2022, ICCV 2021, IEEE FG 2020 and IEEE CVPR 2017, with the participation of many teams coming from both academia and industry, from all across the world [113, 108, 97, 84, 91, 96, 101, 115, 58, 114, 113, 84, 100, 104, 17, 73, 71, 115, 85, 28, 10, 65, 112, 12, 6, 4, 11, 110, 90, 92, 106, 98, 25, 1, 75, 62, 14, 13, 102, 68, 15, 69, 76, 23, 19, 9, 80, 70, 107, 74, 82, 26, 79, 111, 103, 27, 77, 99, 21, 95, 20, 89, 94, 24, 8, 18, 86, 5, 16, 66, 93, 72, 83, 67, 109, 63, 22, 64, 56, 3, 57, 55, 2, 53, 49, 52, 61, 81, 42, 38, 43, 46, 35, 33, 34, 36, 87, 88, 45, 50, 41].
2 Competition Corpora
In the following, we present a brief synopsis of each Challenge’s dataset. For a more comprehensive understanding, readers are encouraged to consult the original documentation. Additionally, we detail the pre-processing steps undertaken for the first three Challenges, which involve cropping and aligning all image-frames. These have been utilized in our baseline experiments.
2.1 Valence-Arousal Estimation Challenge
This Challenge’s dataset comprises $594$ videos, an expansion of the Aff-Wild2 database, annotated in terms of valence and arousal. Notably, sixteen videos feature two subjects, both of whom are annotated. In total, annotations are provided for 2,993,081 frames from $584$ subjects; these annotations have been conducted by four experts using the methodology outlined in [7]. Valence and arousal values are continuous and range in $[-1,1]$. The 2D Valence-Arousal histogram of annotations is visualized in Figure 1.
Figure 1: Valence-Arousal Estimation Challenge: 2D Valence-Arousal Histogram of Annotations in Aff-Wild2
Aff-Wild2 is split into training, validation and testing sets, in a subject independent manner, ensuring each subject appears exclusively in one set. The training, validation and testing sets consist of 356, 76 and 162 videos, respectively.
The Train and Validation data along with their corresponding annotations are provided to the participating teams. The unlabeled test data are provided to the participating teams who will upload their test set predictions to an evaluation server. Participating teams are allowed to submit up to five sets of predictions for this challenge.
2.2 Expression Recognition Challenge
In this Challenge, the dataset consists of 548 videos from Aff-Wild2, annotated for the six basic expressions, neutral state, and an ’other’ category representing non-basic expressions. Seven videos feature two subjects, both of whom are annotated. In total, annotations are provided for 2,624,160 frames from 437 subjects. Annotation is conducted by seven experts on a frame-by-frame basis. Table 1 presents the distribution of expression annotations.
Table 1: Expression Classification Challenge: Number of Annotated Images for each Expression
Expressions No of Images
Neutral 468,069
Anger 36,627
Disgust 24,412
Fear 19,830
Happiness 245,031
Sadness 130,128
Surprise 68,077
Other 512,262
Aff-Wild2 is split into training, validation and testing sets, in a subject independent manner. Aff-Wild2 is split into training, validation and testing sets, in a subject independent manner. The training, validation and testing sets consist of 248, 70 and 230 videos, respectively.
The Train and Validation data along with their corresponding annotations are provided to the participating teams. The unlabeled test data are provided to the participating teams who will upload their test set predictions to an evaluation server. Participating teams are allowed to submit up to five sets of predictions for this challenge.
2.3 Action Unit Detection Challenge
The dataset for this Challenge comprises 542 videos annotated for 12 AUs corresponding to the inner and outer brow raiser, the brow lowerer, the cheek raiser, the lid tightener, the upper lip raiser, the lip corner puller and depressor, the lip tightener and pressor, lips part and jaw drop. The exact utulized AUs along with their corresponding actions can be seen in Table 2. Annotations are provided for 2,627,632 frames from 438 subjects. A semi-automatic annotation procedure, involving both manual and automatic annotations, is employed. Table 2 further details the annotated AUs distribution.
Table 2: Action Unit Detection Challenge: Distribution of AU Annotations in Aff-Wild2
Action Unit # Action
Total Number
of Activated AUs
AU 1 inner brow raiser 301,102
AU 2 outer brow raiser 139,936
AU 4 brow lowerer 386,689
AU 6 cheek raiser 619,775
AU 7 lid tightener 964,312
AU 10 upper lip raiser 854,519
AU 12 lip corner puller 602,835
AU 15 lip corner depressor 63,230
AU 23 lip tightener 78,649
AU 24 lip pressor 61,500
AU 25 lips part 1,596,055
AU 26 jaw drop 206,535
Aff-Wild2 is split into training, validation and testing sets, in a subject independent manner. The training, validation and testing sets consist of 295, 105 and 142 videos, respectively.
The Train and Validation data along with their corresponding annotations are provided to the participating teams. The unlabeled test data are provided to the participating teams who will upload their test set predictions to an evaluation server. Participating teams are allowed to submit up to five sets of predictions for this challenge.
2.4 Compound Expression Recognition Challenge
For this Challenge, a part of C-EXPR-DB database will be used (56 videos in total). C-EXPR-DB is audiovisual (A/V) in-the-wild database and in total consists of 400 videos of around 200K frames; each frame is annotated in terms of 12 compound expressions. For this Challenge, the following 7 compound expressions will be considered: Fearfully Surprised, Happily Surprised, Sadly Surprised, Disgustedly Surprised, Angrily Surprised, Sadly Fearful and Sadly Angry.
Goal of the Challenge and Rules Participants will be provided with a part of C-EXPR-DB database (56 videos in total), which will be unannotated, and will be required to develop their methodologies (supervised/self-supervised, domain adaptation, zero-/few-shot learning etc) for recognising the 7 compound expressions in this unannotated part, in a per-frame basis.
2.5 Emotional Mimicry Intensity Estimation Challenge
For this Challenge, the multimodal Hume-Vidmimic2 dataset is used which consists of more than 15,000 videos totaling over 25 hours. In this dataset, every participant was tasked with imitating a ’seed’ video that showcased an individual displaying a particular emotion. Following the mimicry, they were then asked to assess the emotional intensity of the seed video by selecting from a range of predefined emotional categories. The following emotion dimensions are targeted: ’Admiration’, ’Amusement’, ’Determination’, ’Empathic Pain’, ’Excitement’, and ’Joy’. A normalized score from 0 to 1 is provided as a ground truth value.
2.6 Aff-Wild2 Pre-Processing: Cropped & Cropped-Aligned Images
Initially, all videos are segmented into individual frames, after which they undergo processing using the RetinaFace detector. This step aims to extract face bounding boxes and five facial landmarks for each frame. Subsequently, the images are cropped based on the bounding box coordinates, and these cropped images are provided to the participating teams.
Using the five facial landmarks (representing two eyes, the nose, and two mouth corners), a similarity transformation is applied. This transformation ensures alignment, resulting in cropped and aligned images, which are also shared with the participating teams. Ultimately, these cropped and aligned images are utilized in our baseline experiments.
All cropped and cropped-aligned images are resized to dimensions of $112\times 112\times 3$ pixels and their intensity values are normalized to fall within the range of $[-1,1]$.
3 Evaluation Metrics Per Challenge
3.1 Valence-Arousal Estimation Challenge
The performance measure is the average between the Concordance Correlation Coefficient (CCC) of valence and arousal:
$\mathcal{P}_{VA}=\frac{CCC_{a}+CCC_{v}}{2}$ (1)
CCC evaluates the agreement between two time series (e.g., all video annotations and predictions) by scaling their correlation coefficient with their mean square difference. In this way, predictions that are well correlated with the annotations but shifted in value are penalized in proportion to the deviation. CCC takes values in the range $[-1,1]$, where $+1$ indicates perfect concordance and $-1$ denotes perfect discordance. The highest the value of the CCC the better the fit between annotations and predictions, and therefore high values are desired. CCC is defined as follows:
$CCC=\frac{2s_{x}s_{y}\rho_{xy}}{s_{x}^{2}+s_{y}^{2}+(\bar{x}-\bar{y})^{2}},$ (2)
where $\rho_{xy}$ is the Pearson’s Correlation Coefficient, $s_{x}$ and $s_{y}$ are the variances of all video valence/arousal annotations and predicted values, respectively and $s_{xy}$ is the corresponding covariance value.
3.2 Expression Classification Challenge
The performance measure is the average F1 Score across all 8 categories (i.e., macro F1 Score):
$\mathcal{P}_{EXPR}=\frac{\sum_{expr}F_{1}^{expr}}{8}$ (3)
The $F_{1}$ score is a weighted average of the recall (i.e., the ability of the classifier to find all the positive samples) and precision (i.e., the ability of the classifier not to label as positive a sample that is negative). The $F_{1}$ score takes values in the range $[0,1]$; high values are desired. The $F_{1}$ score is defined as:
$F_{1}=\frac{2\times precision\times recall}{precision+recall}$ (4)
3.3 Action Unit Detection Challenge
The performance measure is the average F1 Score across all 12 AUs. Therefore, the evaluation criterion for the Action Unit Detection Challenge is:
$\mathcal{P}_{AU}=\frac{\sum_{au}F_{1}^{au}}{12}$ (5)
3.4 Compound Expression Recognition Challenge
The performance measure is the average F1 Score across all 7 compound expressions. Therefore, the evaluation criterion for the Compound Expression Recognition Challenge is:
$\mathcal{P}_{CE}=\frac{\sum_{expr}F_{1}^{expr}}{7}$ (6)
3.5 Emotional Mimicry Intensity Estimation Challenge
The performance measure is the average Pearson’s Correlation Coefficient ($\rho$) across the 6 emotion dimensions:
$\mathcal{P}_{EMI}=\frac{\sum_{i=1}^{6}\rho^{i}}{6}$ (7)
Pearson’s Correlation Coefficient ($\rho$) takes values in the range $[-1,1]$; high values are desired.
4 Participating Teams’ and Baseline Methods’ Results
All baseline systems are built solely on existing open-source machine learning toolkits to maintain result reproducibility. TensorFlow is the chosen framework for implementing all systems.
In this Section, we describe the baseline systems developed for each Challenge and present their obtained results.
4.1 Valence-Arousal Estimation Challenge
The baseline network comprises a ResNet architecture with 50 layers, initially trained on ImageNet (ResNet50). It incorporates a linear output layer responsible for providing the final estimations for valence and arousal. Its average CCC performance on the validation set is 0.22 (the CCC for valence is 0.24 and the CCC for arousal is 0.20), as can be seen in Table 3.
Table 3: Valence-Arousal Estimation Challenge’s Results; Total Score is the average CCC between valence and arousal
Teams
Total Score
CCC-V
CCC-A
baseline 0.22
0.24
0.20
4.2 Expression Classification Challenge
The baseline network adopts a VGG16 architecture with fixed convolutional weights (i.e., non-trainable), while only the three fully connected layers are trainable. It is pre-trained on the VGGFACE dataset and equipped with an output layer featuring a softmax activation function, facilitating the prediction of eight expressions. Mixaugment has been used as data augmentation technique.
MixAugment is a simple and data-agnostic data augmentation routine that trains a method on convex combinations of pairs of examples and their labels. It extends the training distribution by incorporating the prior knowledge that linear interpolations of feature vectors should lead to linear interpolations of the associated targets. MixAugment constructs virtual training examples $(\tilde{x},\tilde{y})$ as follows:
$\tilde{x}$ $=\lambda x_{i}+(1-\lambda)x_{j}$
$\tilde{y}$ $=\lambda y_{i}+(1-\lambda)y_{j}$ (8)
where $x_{i}$ and $x_{j}$ are two random raw inputs (i.e., images), $y_{i}$ and $y_{j}$$\in\{0,1\}^{8}$ are their corresponding one-hot label encodings and $\lambda\thicksim\mathrm{B}(\alpha,\alpha)\in[0,1]$ (i.e., Beta distribution) for $\alpha\in(0,\infty)$.
During each training iteration, the baseline network is trained concurrently on both real (r) and virtual (v) examples. Specifically, in each training iteration, the network is fed with both $x_{i}$ and $x_{j}$, and the generated image $\tilde{x}$ (of Eq. 8).
The performance of the baseline network on the validation set is 0.25, as can be seen in Table 6.
Table 4: Expression Recognition Challenge’s Results; the evaluation criterion is the average F1 Score
Teams
Total Score
baseline 0.25
4.3 Action Unit Detection Challenge
The baseline network adopts a VGG16 architecture with fixed convolutional weights (i.e., non-trainable), while only the three fully connected layers are trainable. It is pre-trained on the VGGFACE dataset and equipped with an output layer featuring a sigmoid activation function, facilitating the detection of the twelve AUs. Its performance on the validation set is 0.39, as can be seen in Table 6.
Table 5: Action Unit Detection Challenge’s Results; the evaluation criterion is the average F1 Score
Teams
Total Score
baseline 0.39
4.4 Emotional Mimicry Intensity Estimation Challenge
We established baseline results using two different feature sets.
First, we employed pre-trained Vision Transformer (ViT) features, which were further processed through a three-layer Gated Recurrent Unit (GRU) network. This approach achieved a performance score of 0.09.
Secondly, we utilized features extracted from Wav2Vec2, combined with a linear processing layer, which resulted in a performance score of 0.24.
Additionally, we explored a multimodal approach by averaging the predictions from both unimodal methods, leading to a combined performance score of 0.48.
Table 6: Emotional Mimicry Intensity Estimation Challenge’s Results; the evaluation criterion is the average Pearson’s Correlation
Teams
Total Score
ViT-GRU baseline 0.09
Wav2Vec2-FC 0.24
ViT-GRU + Wav2Vec2-FC 0.48
5 Conclusion
In this paper we have presented the sixth Affective Behavior Analysis in-the-wild Competition (ABAW) held in conjunction with IEEE CVPR 2024. This Competition is a continuation of the series of ABAW Competitions. This Competition comprises five Challenges targeting: i) Valence-Arousal Estimation, ii) Expression Recognition (8 categories), iii) Action Unit Detection (12 action units), iv) Compound Expression Recognition (7 categories) and v) Emotional Mimicry Intensity Estimation (6 emotion dimensions). The databases utilized for this Competition are an extended version of Aff-Wild2, the C-EXPR-DB and the Hume-Vidmimic2 dataset.